OpenStudy (anonymous):

Fun Land theme park has a total of 110 males and females who are employed at the park. If the number of males is 10 more than three times the number of females, how many males are employed at the park?

OpenStudy (anonymous):

We need two equations, can you find one of them, then I will find the other?

OpenStudy (anonymous):

There are two equations: f+m=110 and 3f+10=m. Substitute m in f+m for 3f+10 to get f+3f+10=110 add the fs and subtract the 10 to get 4f=100 f=25, so there were 25 females. You can find the amount of males.

OpenStudy (anonymous):

do you understand?